# Understanding Oversmoothing in GNNs as Consensus in Opinion Dynamics

### v1

@@ 4.1 Oversmoothing is Opinion Consensus
4.2 Oversmoothing in GRAND
Using our graph neural network, opinion dynamics analogy we show that the recently introduced continuous-depth GNN model graph neural diffusion (GRAND) (Chamberlain et al., 2021b) and its variants exhibit oversmoothing in the limit.
Linear GRAND ($-\ell$).
In GRAND, information is exchanged between nodes according to the diffusion process,
$\frac{\partial\mathbf{X}(t)}{\partial t}=\operatorname{div}\left[\mathbf{G}(\mathbf{X}(t),t)\,\nabla\mathbf{X}(t)\right],$ (7)
where $\mathbf{G}=\operatorname{diag}(a(\mathbf{x}_{i},\mathbf{x}_{j},t))$, for some similarity measure $a(\cdot)$ evaluated over all connected nodes (i.e., $e_{ij}\in\mathcal{E}$).
Linear GRAND ($-\ell$), is a special case of GRAND with time-independent $\mathbf{G}$. In this case, the similarity measure is defined using multi-head attention (Vaswani, 2017),
@@ 4.2 Oversmoothing in GRAND
$a(\mathbf{x}_{i},\mathbf{x}_{j})=\operatorname{softmax}\left(\frac{(\mathbf{W}_{K}\mathbf{x}_{i})^{\top}\mathbf{W}_{Q}\mathbf{x}_{j}}{d_{k}}\right),$ (8)
where $\mathbf{W}_{K}$ and $\mathbf{W}_{Q}$ are learned matrices with dimension $d_{k}\times n$. In this case, the dynamics can be expressed,
$\dot{\mathbf{X}}(t)=(\mathbf{A}-\mathbf{I})\mathbf{X}(t),$ (9)
where $\mathbf{A}=a(\mathbf{x}_{i},\mathbf{x}_{j})$. Since $\mathbf{A}$, is right stochastic, the dynamical equation for $-\ell$ is equivalent to the dynamical equation for linear opinion dynamics (see Equation 5) and in the limit, $-\ell$ will exhibit oversmoothing.
Proposition 4.1.
In the limit GRAND-$l$ will exhibit oversmoothing.
@@
By distributing and rearranging terms in Equation(9), the dynamics of $-\ell$ can be expressed,
$\mathbf{\dot{X}}(t)=-\mathbf{X}(t)+\mathbf{A}\mathbf{X}(t).$ (10)
Since $\mathbf{A}$ is constrained to be right stochastic the dynamics are equivalent to those of the linear opinion dynamics model in Equation (4). Consequently, all nodes will converge to the same value. This is the consensus/oversmoothing condition described in Equations (6) and (2). ∎
Adding a damping term does not resolve oversmoothing.
GraphCON (Rusch et al., 2022) introduces a framework to mitigate oversmoothing in continuous-depth GNNs by stabilizing the time evolution of the Dirichlet energy. This is achieved by amending an arbitrary GNN with damping terms and converting it to a continuous-depth model. The dynamical update for a GraphCON amended GNN $F_{\theta}$ is given by,
$\dot{\mathbf{Y}}(t)=\sigma(F_{\theta}(\mathbf{X}(t),t))-\gamma\mathbf{X}(t)-\alpha\mathbf{Y}(t),\quad\dot{\mathbf{X}}(t)=\mathbf{Y}(t),$ (11)
where $\sigma$ is the activation function, and $\gamma$ and $\alpha$ are damping terms that force the dynamics converge to a desired equilibrium.
In GraphCON-Tran (Rusch et al., 2022), the authors amend $-\ell$ with the GraphCON framework and achieve improved performance at increased network depths. The dynamics of the GraphCON-Tran model is given by
$\dot{\mathbf{Y}}(t)=\mathbf{(A-I)X}(t)-\mathbf{Y}(t),\quad\dot{\mathbf{X}}(t)=\mathbf{Y}(t).$ (12)
Here $\sigma$ is the identity mapping, $\gamma=1$, and $F_{\theta}(\mathbf{X},t)=\mathbf{AX}(t)$. The corresponding steady solution is
$\mathbf{0}=\mathbf{(A-I)X}(T),\quad\mathbf{0}=\mathbf{Y}(T),$ (13)
which is equivalent to $-\ell$ (see Equation (9)), so GraphCON-Tran will exhibit oversmoothing.
Adding a source term is not the perfect answer.
In GRAND++ (Thorpe et al., 2022), the authors amend GRAND with a source term to alleviate oversmoothing. The dynamical update for GRAND++ is given by,
$\frac{\partial\mathbf{x}(t)^{(i)}}{\partial t}=\operatorname{div}\left[\mathbf{G}(\mathbf{X}(t),t)\,\nabla\mathbf{X}(t)\right]^{(i)}+\sum_{j\in\mathcal{I}}\delta_{ij}\mathbf{b}_{j},$ (14)
where $\mathcal{I}\subseteq\{1,\dots,n\}$ is the index set of nodes whose features are used to compute the source term, and $\mathbf{b}_{j}$ is a function of $\mathbf{X}(0)$.
$-\ell$ is a special case of GRAND++ with time-independent $\mathbf{G}$,
$\dot{\mathbf{X}}(t)=(\mathbf{A}-\mathbf{I})\mathbf{X}(t)+\mathbf{B}.$ (15)
While the performance of $-\ell$ exceeds that of $-\ell$, the performance boost comes at the expense of the system equilibrium. Moreover, for large timescales, the difference in node features is relatively small resulting in reduced discriminability, and poorer network performance.
Proposition 4.2.
As $t\to\infty$ the dynamics of $-\ell$ are approximately linear in $t$.
We prove Proposition 4.2 in Appendix A.1.
5 BIMP: Behavior-Inspired Message Passing Neural Network
In this section, we introduce our BIMP model. We show that by incorporating nonlinear opinion dynamics as an inductive bias we circumvent oversmoothing and retain expressivity.
5.1 Formulation
In our proposed BIMP model, information is exchanged between nodes according the nonlinear opinion dynamics model,
$\dot{\mathbf{X}}=-d\mathbf{X}+\tanh\bigg{[}u\bigg{(}\alpha\mathbf{X}+\mathbf{A}^{\mathrm{a}}\mathbf{X}\\ +\mathbf{X}{\mathbf{A}^{\mathrm{o}}}^{\top}+\mathbf{A}^{\mathrm{a}}\mathbf{X}{\mathbf{A}^{\mathrm{o}}}^{\top}\bigg{)}\bigg{]}+\mathbf{B},$ (16)
where we set $d_{ij}=d$, $\alpha_{ij}=\alpha$ and $u_{i}=u$. The flexibility of nonlinear opinion dynamics allows BIMP to model more diverse node feature representations than approaches whose dynamics are equivalent to linear opinion dynamics.
5.2 $-\ell$ is a first-order approximation of BIMP
Proposition 5.1.
For a BIMP model with attention parameter $u=1$, bias parameter $\mathbf{B=0}$, and uncorrelated options $\mathbf{A^{\mathrm{o}}}=\mathbf{0}$, the first-order approximation is equivalent to $-\ell$.
@@ 5.2 $-\ell$ is a first-order approximation of BIMP
The BIMP model described in the proposition has dynamics of the form
When $c=1$, this equation reduces to,
$\dot{\mathbf{X}}=(\mathbf{A}^{\mathrm{a}}-\mathbf{I})\mathbf{X},$ (19)
which is of the same form as $-\ell$ (see Equation (7)). Since $-\ell$ has the same form as the first-order approximation of our BIMP model, we say our BIMP model has greater expressive capacity. ∎
5.3 Designing an expressive BIMP model
The effective adjacency matrix.
To ensure our model has the expected behavior, the effective adjacency matrix must be constrained to have a simple leading eigenvalue (Leonard et al., 2024). In (Leonard et al., 2024) the authors consider a nonlinear opinion dynamics model of the form
where the communication matrix $\mathbf{A}^{\mathrm{a}}$ has a simple leading eigenvalue. We consider the general form of nonlinear opinion dynamics given in Equation (16) and show that it can be reduced to Equation (20) with the following definition for the effective adjacency matrix.
Definition 5.2.
Given the communication matrix $\mathbf{A}^{\mathrm{a}}$ and belief matrix $\mathbf{A}^{\mathrm{o}}$, the effective adjacency matrix $\tilde{\mathbf{A}}$ is defined
$\tilde{\mathbf{A}}=(\mathbf{A}^{\mathrm{o}}+\mathbf{I})\otimes(\mathbf{A}^{\mathrm{a}}+\mathbf{I}),$ (21)
where $\otimes$ denotes the Kronecker product.
Given this definition, we have the following proposition.
Proposition 5.3.
Using the effective adjacency matrix, the general form of nonlinear opinion dynamics (Equation (16)) can be reduced to the form where options are uncorrelated (Equation (20)).
@@ 5.3 Designing an expressive BIMP model
We first rewrite Equation (16) in the following form:
$\dot{\mathbf{X}}=-d\mathbf{X}+\tanh\Big{[}u\Big{(}(\alpha-1)\mathbf{X}\\ +(\mathbf{A}^{\mathrm{a}}+\mathbf{I})\mathbf{X}({\mathbf{A}^{\mathrm{o}}}^{\top}+\mathbf{I})\Big{)}\Big{]}+\mathbf{B}.$ (22)
From here, we can write the matrix product $\mathbf{ABC}$ with $\mathbf{A}\in\mathbb{R}^{m\times m}$, $\mathbf{B}\in\mathbb{R}^{m\times n}$ and $\mathbf{C}\in\mathbb{R}^{n\times n}$, as $\mathbf{ABC}=\operatorname{vec}^{-1}\left[\left(\mathbf{C}^{\top}\otimes\mathbf{A}\right)\operatorname{vec}(\mathbf{B})\right]$, where $\operatorname{vec}$ denotes the vectorization operator (Magnus & Neudecker, 2019). This yields the following form
$\dot{\mathbf{X}}=-d\mathbf{X}+\tanh\left[u\left(\tilde{\alpha}\mathbf{X}+\operatorname{vec^{-1}}(\tilde{\mathbf{A}}\operatorname{vec}(\mathbf{X})\right)\right]+\mathbf{B},$ (23)
where $\tilde{\alpha}=\alpha-1$, and $\tilde{\mathbf{A}}=(\mathbf{A}^{\mathrm{o}}+\mathbf{I})\otimes(\mathbf{A}^{\mathrm{a}}+\mathbf{I})$ follows from Definition 5.2. Vectorizing both sides of yields
$\dot{\mathbf{x}}=-d\mathbf{x}+\tanh\left[u\left(\tilde{\alpha}\mathbf{x}+\tilde{\mathbf{A}}\mathbf{x}\right)\right]+\mathbf{b},$ (24)
where $\mathbf{x}=\operatorname{vec}(\mathbf{X})$, and $\mathbf{b}=\operatorname{vec}(\mathbf{B})$. We obtain the general nonlinear opinion dynamics in the form (20). ∎
Leveraging Proposition 5.3 we can say that if the effective adjacency matrix has a simple leading eigenvalue the system will exhibit the expected behavior.
The condition on the effective adjacency matrix is satisfied by a right-stochastic matrix (the leading eigenvalue of a right-stochastic matrix is $\lambda_{\max}=1$), however, the size of the effective adjacency matrix (i.e., $\mathbb{R}^{N_{\mathrm{o}}N_{\mathrm{a}}\times N_{\mathrm{o}}N_{\mathrm{a}}}$) would impose a computational challenge. To circumvent this, we instead learn the communication matrix and belief matrix separately. Specifically, the entries of the communication matrix $\mathbf{A}^{a}=[a^{\mathrm{a}}_{ik}]$ are defined using multi-head attention,
$a^{\mathrm{a}}_{ik}=\operatorname{softmax}\left(\frac{(\mathbf{W}^{\mathrm{a}}_{K}\mathbf{x}_{i})^{\top}\mathbf{W}^{\mathrm{a}}_{Q}\mathbf{x}_{k}}{d_{k}^{\mathrm{a}}}\right),$ (25)
where $\mathbf{W}^{\mathrm{a}}_{K}$ and $\mathbf{W}^{\mathrm{a}}_{Q}$ are the key and query weight matrices, and $d_{k}^{\mathrm{a}}$ is the attention dimension. $\mathbf{x}_{i}$ is the opinion of agent $i$. The entries of the belief matrix $\mathbf{A}^{o}=[a^{\mathrm{o}}_{jl}]$ are defined similarly,
$a^{\mathrm{o}}_{jl}=\operatorname{softmax}\left(\frac{(\mathbf{W}^{\mathrm{o}}_{K}\mathbf{x}^{\top}_{j})^{\top}\mathbf{W}^{\mathrm{o}}_{Q}\mathbf{x}^{\top}_{l}}{d_{k}^{\mathrm{o}}}\right),$ (26)
where $\mathbf{W}^{\mathrm{o}}_{K}$ and $\mathbf{W}^{\mathrm{o}}_{Q}$ are the key and query weight matrices, and $d_{k}^{\mathrm{o}}$ is the attention dimension.
The attention parameter.
For our model to utilize its expressive capacity, it must not degenerate to a linear model. The degenerate form occurs both when $u$ is very small, since the nonlinear term evaluates to zero, and when $u$ is very large, since the nonlinear term evaluates to one. To avoid this, we set the attention to the critical point of the attention-opinion bifurcation diagram.
Proposition 5.4.
The critical point $u$ of the attention-opinion bifurcation diagram is equal to $\nicefrac{{d}}{{\tilde{\alpha}+\lambda^{\tilde{a}}_{\max}}}$.

### v2

@@ 3.2 Opinion Dynamics and Opinion Consensus
3.3 Graph Neural Network-Opinion Dynamics Analogy
As described in Section 3.1 and 3.2, GNNs and opinion dynamics models have several features which can be understood analogously. The nodes in a GNNs are analogous to the agents in an opinion dynamics model, the edges between nodes are analogous to communication links between agents, and layer-wise oversmoothing is analogous to opinion consensus in time (see Equations (3) and (4)). We summarize our GNN-opinion dynamics (GNN-OD) analogy in Table 1.
4 Oversmoothing and Opinion Consensus
In this section, we use our analogy to prove oversmoothing in several classes of GNNs. We begin with linear discrete-depth GNNs, then continue to Laplacian-based continuous-depth GNNs. The utility of the GNN-OD analogy for understanding oversmoothing, motivates its use in the design of new architectures (Section 5). All proofs are provided in Appendix A.
4.1 Linear discrete-depth GNNs
Linear discrete-depth GNNs (e.g., SGC [76] and DGC [77]) can be described using layer-wise transformations of the form,
$\mathbf{X}^{l+1}=\mathbf{AX}^{l}\mathbf{W}^{l}=\mathbf{D}\tilde{\mathbf{A}}\mathbf{X}^{l}\mathbf{W}^{l}=\tilde{\mathbf{A}}\mathbf{X}^{l}\mathbf{D}\mathbf{W}^{l},$ (8)
where adjacency matrix $\mathbf{A}=\mathbf{D}\tilde{\mathbf{A}}$ for some right stochastic matrix $\tilde{\mathbf{A}}$ and diagonal matrix $\mathbf{D}$ with $\mathbf{D}_{ii}=\sum_{j}\mathbf{A}_{ij}$; and the transformation matrix $\mathbf{W}^{l}$ is layer dependent. We can use our GNN-OD analogy to show this class of GNNs will exhibit oversmoothing for all inputs, and all input graphs.
Lemma 4.1 (Linear dynamics oversmooth).
Any discrete-depth graph neural network with linear aggregation exhibits oversmoothing.
Previous works that have shown oversmoothing in various subclasses of linear discrete-depth GNNs include [39] which proves oversmoothing in convolutional GNNs, and [46] which proves oversmoothing in discrete-depth GNNs with linear mean aggregation.
4.2 Laplacian-based continuous-depth GNNs
Laplacian dynamics. For a given graph, the graph Laplacian of is defined as $\mathbf{L=A-D}$ where $\mathbf{A}$ is the graph adjacency matrix and $\mathbf{D}$ is a diagonal matrix with entries $\mathbf{D}_{ii}=\sum_{j}\mathbf{A}_{ij}$. GNNs with Laplacian dynamics (e.g., $-\ell$ [26] and GraphCON-Tran [32]), can be described using the dynamical equation,
$\dot{\mathbf{X}}(t)=-\mathbf{LX}(t).$ (9)
We can use our GNN-OD analogy to show this class of GNNs will exhibit oversmoothing for all inputs, and all input graphs.
Lemma 4.2 (Laplacian dynamics oversmooth).
Any continuous-depth graph neural network with Laplacian dynamics exhibits oversmoothing.
Oversmoothing of continuous-depth GNNs with Laplacian dynamics and time-varying adjacency matrix (e.g., GRAND-$n\ell$ [26]), can be shown by analogy with linear opinion dynamics models with time-varying influence (see Section 3.2). Oversmoothing in $-\ell$ [26] was previously shown in [44, 48, 49].
Laplacian dynamics with an external input. Continuous-depth GNNs with Laplacian dynamics and an external input (e.g., [33, 48, 49, 78, 79]), can be described using the dynamical equation,
$\dot{\mathbf{X}}(t)=-\mathbf{L}\mathbf{X}(t)+\mathbf{B}(t),$ (10)
where $\mathbf{B}(t)$ is the external input. While some works [33, 48, 49] design $\mathbf{B}(t)$ to address oversmoothing, using our GNN-OD analogy, we can understand when oversmoothing occurs.
Lemma 4.3 (Laplacian dynamics with an external input oversmooth).
Any continuous-depth graph neural network with Laplacian dynamics and an external input will exhibit oversmoothing when the dynamics are linear.
5 Behavior-Inspired Message Passing Neural Network
In this section, we describe our Behavior-Inspired Message Passing (BIMP) GNN which leverages the nonlinear opinion dynamics model as an inductive bias. The nonlinear opinion dynamics model is more general than the linear opinion dynamics model, and can be designed to converge to dissensus for general inputs. We begin with our model definition, then prove several desirable properties including robustness to oversmoothing, well behaved gradients, and adaptability to homophilic and heterophilic datasets. All proofs are provided in Appendix A.
5.1 Model definition
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be an input graph with $n$ nodes, where each node has a $d_{\text{in}}$-dimensional feature representation. BIMP applies a learnable encoder $\phi$, decoder $\psi$, and nonlinear opinion dynamics model to produce an output. The encoder is defined $\phi:\mathbb{R}^{d_{\text{in}}}\to\mathbb{R}^{N_{o}}$, the decoder is defined $\psi:\mathbb{R}^{N_{\mathrm{o}}}\to\mathbb{R}^{d_{\text{class}}}$, and our dynamics are defined by the equation,
In our dynamics, the parameters $d$ and $\alpha$ are hyperparameters, the attention parameter $u=\frac{d}{\alpha+3}$, the initial condition $\mathbf{X}(0)=\phi(\mathbf{X}_{\text{in}})$, and the adjacency matrices of communication and option graphs, $\mathbf{A}^{\mathrm{a}}$ and $\mathbf{A}^{\mathrm{o}}$ respectively, are learned. The output of our model is given by $\mathbf{Y}=\psi(\mathbf{X}(T))$, where the terminal time $T$ is a hyperparameter.
5.1.1 The communication and option graphs.
The communication and option graphs are designed to allow for theoretical analysis, and reduce computational expense. For a nonlinear opinion dynamics model of the form,
the communication adjacency matrix $\mathbf{A}^{\mathrm{a}}$ must have a simple leading eigenvalue to admit analysis [35]. Our BIMP model described in Equation (11) can be written in the form of Equation (12) by combining the communication and option graphs into a single effective adjacency matrix.
Definition 5.1 (Effective adjacency matrix).
Given the adjacency matrices of the communication and option graphs, $\mathbf{A}^{\mathrm{a}}$ and $\mathbf{A}^{\mathrm{o}}$, the effective adjacency matrix $\tilde{\mathbf{A}}$ is defined, $\tilde{\mathbf{A}}=(\mathbf{A}^{\mathrm{o}}+\mathbf{I})\otimes(\mathbf{A}^{\mathrm{a}}+\mathbf{I}).$ (13)
Lemma 5.2.
The general form of nonlinear opinion dynamics (Equation (11)) can be written, $\dot{\mathbf{x}}=-d\mathbf{x}+\tanh\left[u\biggl{(}(\alpha-1)\mathbf{x}+\tilde{\mathbf{A}}\mathbf{x}\biggr{)}\right]+\mathbf{b},$ (14) where $\tilde{\mathbf{A}}$ is the effective adjacency matrix, $\mathbf{x}=\operatorname{vec}(\mathbf{X})$ and $\mathbf{b}=\operatorname{vec}(\mathbf{B})$.
In order to analyze the behavior of BIMP, the effective adjacency matrix must be constrained to have a simple leading eigenvalue. We enforce this condition by constraining the learned effective adjacency matrix to be right stochastic (the leading eigenvalue of a right stochastic matrix is $\lambda_{\max}=1$). Learning this matrix directly would be computationally prohibitive (the size of the effective adjacency matrix is $\mathbb{R}^{N_{\mathrm{o}}N_{\mathrm{a}}\times N_{\mathrm{o}}N_{\mathrm{a}}}$). To relieve the computational burden, we instead learn the communication and option graphs separately. Specifically, the entries of $\mathbf{A}^{\mathrm{a}}=[a^{\mathrm{a}}_{ik}]$ and $\mathbf{A}^{\mathrm{o}}=[a^{\mathrm{o}}_{jl}]$ are defined using multi-head attention,
@@ 5.1 Model definition
$a^{\mathrm{a}}_{ik}=\operatorname{softmax}\left(\frac{(\mathbf{W}^{\mathrm{a}}_{K}\mathbf{x}_{i})^{\top}\mathbf{W}^{\mathrm{a}}_{Q}\mathbf{x}_{k}}{d_{k}^{\mathrm{a}}}\right),\quad a^{\mathrm{o}}_{jl}=\operatorname{softmax}\left(\frac{(\mathbf{W}^{\mathrm{o}}_{K}\mathbf{x}^{\top}_{j})^{\top}\mathbf{W}^{\mathrm{o}}_{Q}\mathbf{x}^{\top}_{l}}{d_{k}^{\mathrm{o}}}\right),$ (15)
where $\mathbf{W}^{\mathrm{a}}_{K}$, $\mathbf{W}^{\mathrm{a}}_{Q}$, $\mathbf{W}^{\mathrm{o}}_{K}$ and $\mathbf{W}^{\mathrm{o}}_{Q}$ are the key and query weight matrices for communication and option graphs. A useful consequence of this approach is that the leading eigenvalue $\lambda^{\tilde{a}}_{\max}$ of $\tilde{\mathbf{A}}$ is constant and does not need to be recomputed during training.
Lemma 5.3.
The leading eigenvalue of the effective adjacency matrix $\tilde{\mathbf{A}}$ is $\lambda^{\tilde{a}}_{\max}=4$.
5.2 Parameter selection
The attention parameter $u$. In the nonlinear opinion dynamics model, the attention parameter $u$ is the bifurcation parameter. Near the bifurcation point $u^{*}$ (i.e., the point where the number and/or stability of solutions change), the model is ultrasensitive to the input $\mathbf{B}$, and agents will quickly form an opinion (see Figure 1). We design BIMP to be ultrasensitive to the input by setting the value of the attention parameter $u$ to the bifurcation point of the attention-opinion bifurcation diagram.
Lemma 5.4 (Bifurcation point $u^{*}$).
The bifurcation point $u^{*}$ of the attention-opinion bifurcation diagram is equal to $\nicefrac{{d}}{{\left(\alpha-1+\lambda^{\tilde{a}}_{\max}\right)}}$.
From Lemma 5.4 and 5.3, the value of the attention parameter at the bifurcation point is $u=\frac{d}{\alpha+3}$.
The input parameter $\mathbf{B}$. In the nonlinear opinion dynamics model, the input parameter $\mathbf{B}$ transforms the bifurcation diagram from a symmetric pitchfork bifurcation to an unfolded pitchfork bifurcation (see Figure 1). This is a form of selective ultrasensitivity where the directions of ultrasensitivity are determined by the structure of the communication graph [36, 35].
In BIMP with effective adjacency matrix $\tilde{\mathbf{A}}$ and attention parameter $u=\frac{d}{\alpha+3}$, oversmoothing depends on the choice of input parameter. We select the input parameter $\mathbf{B}$, so that BIMP converges to an equilibrium for all initial opinions.
Lemma 5.5 (BIMP converges to equilibrium).
A BIMP model with a constant input parameter $\mathbf{B}$ converges to an equilibrium.
A consequence of this design choice is that we can now understand the oversmoothing behavior of BIMP by analyzing its equilibrium behavior. When the input parameter $\mathbf{B}$ is nonzero, BIMP will not exhibit oversmoothing.
Theorem 5.6 (Dissensus in BIMP).
BIMP will not exhibit oversmoothing when the input parameter $\mathbf{B}$ is time-independent with unique entries.
To ensure $\mathbf{B}$ satisfies the conditions of Theorem 5.6, we set $\mathbf{B}=\mathbf{X}(0)$. A similar strategy is used GRAND++ [48], GREAD [49] and KuramotoGNN [50].
5.3 Emergent properties
Table 2: Classification accuracy on homophilic datasets. Classification accuracy on the Cora, citeseer, Pubmed, CoauthorCS, Computers, and Photo datasets are reported. Our BIMP model outperforms competitive baselines on all datasets. We highlight the best and second best accuracy.
Dataset Cora citeseer Pubmed CoauthorCS Computers Photo
BIMP 83.19$\pm$1.13 71.09$\pm$1.40 80.16$\pm$2.03 92.48$\pm$0.26 84.73$\pm$0.61 92.90$\pm$0.44
$-\ell$ 82.20$\pm$1.45 69.89$\pm$1.48 78.19$\pm$1.88 90.23$\pm$0.91 82.93$\pm$0.56 91.93$\pm$0.39
$-\ell$ 82.83$\pm$1.31 70.26$\pm$1.46 78.89$\pm$1.96 90.10$\pm$0.78 82.79$\pm$0.54 91.51$\pm$0.41
KuramotoGNN 81.16$\pm$1.61 70.40$\pm$1.02 78.69$\pm$1.91 91.05$\pm$0.56 80.06$\pm$1.60 92.77$\pm$0.42
GraphCON-Tran 82.80$\pm$1.34 69.60$\pm$1.16 78.85$\pm$1.53 90.30$\pm$0.74 82.76$\pm$0.58 91.78$\pm$0.50
GAT 79.76$\pm$1.50 67.70$\pm$1.63 76.88$\pm$2.08 89.51$\pm$0.54 81.73$\pm$1.89 89.12$\pm$1.60
GCN 80.76$\pm$2.04 67.54$\pm$1.98 77.04$\pm$1.78 90.98$\pm$0.42 82.02$\pm$1.87 90.37$\pm$1.38
GraphSAGE 79.37$\pm$1.70 67.31$\pm$1.63 75.52$\pm$2.19 90.62$\pm$0.42 76.42$\pm$7.60 88.71$\pm$2.68
Table 3: Classification accuracy on heterophilic datasets. Classification accuracy on the Texas, Wisconsin, and Cornell datasets are reported, where BIMP outperforms competitive baselines.
Dataset Cornell Texas Wisconsin
Homophily level 0.30 0.11 0.21
BIMP 77.13$\pm$3.38 82.16$\pm$4.06 86.57$\pm$4.33
$-\ell$ 70.00$\pm$6.22 74.59$\pm$5.43 82.75$\pm$3.90
$-\ell$ 70.30$\pm$8.50 76.14$\pm$5.77 83.09$\pm$2.83
KuramotoGNN 76.02 $\pm$2.77 81.81$\pm$4.36 85.09$\pm$4.42
GraphCON-GCN 74.05$\pm$3.24 80.54$\pm$4.49 84.79$\pm$2.51
GAT 42.16$\pm$7.07 57.84$\pm$5.82 49.61$\pm$4.21
GCN 41.35$\pm$4.69 57.03$\pm$5.98 48.43$\pm$5.75
GraphSAGE 70.54$\pm$2.55 72.70$\pm$5.47 73.14$\pm$6.27
Well behaved gradients. The stability of neural network gradients impacts training efficiency and learning outcomes [32, 50, 80, 81, 82]. Exploding gradients make learning unstable and vanishing gradients slow down learning. In GNNs, oversmoothing has been shown to occur when network gradients vanish [32]. In BIMP, we find that the structure of the nonlinear opinion dynamics inductive bias yields bounded gradients that do not vanish exponentially, even for very deep architectures.
Theorem 5.7 (BIMP has well behaved gradients).
BIMP gradients are upper bounded and do not vanish exponentially.
Model adaptability. In many GNNs, neighborhood aggregation can be interpreted as low-pass filtering [11, 83, 84]. A direct consequence is that these same GNNs will perform poorly on heterophilic datasets (i.e., datasets where edges in an input graphs connect dissimilar nodes). To address this issue, previous works have incorporated high-pass filters which have a sharpening effect [34, 44, 49]. In BIMP, we find that the nonlinear opinion dynamics inductive bias can be interpreted as a tunable filter.
This becomes clear by writing the BIMP dynamics, defined in Equation (14), in the alternative form,
$\dot{\mathbf{x}}=-d\mathbf{x}+\tanh\Bigl{[}u\bigl{(}(\alpha-1)\underbrace{(\mathbf{x-\tilde{A}x})}_{\mathclap{\text{sharpening}}}+\alpha\underbrace{\vphantom{(}\mathbf{\tilde{A}x}}_{\mathclap{\text{smoothing}}}\bigr{)}\Bigr{]}+\mathbf{b}.$ (16)
In this form, BIMP has a high pass filter when $\alpha>1$ ($(\mathbf{x-\tilde{A}x)}$ sharpens the features); and only a low-pass filter when $\alpha\leq 1$ ($\mathbf{\tilde{A}x}$ smooths the feature). By tuning the hyperparameter $\alpha$, BIMP can be adapted to both homophilic and heterophilic datasets.
6 Empirical Analysis
Figure 2: Classification accuracy and Dirichlet energy. BIMP is designed to learn node representations that resist oversmoothing even for very large depths. (Left) We compare the classification accuracy of BIMP to baseline models for architectures with $1,2,4,8,16,32,64$ and $128$ timesteps. Our BIMP model is stable out to 128 timesteps, while baseline performance deteriorates after 32 timesteps. (Right) We compare the Dirichlet energy of node features over a range of network depths. The Dirichlet energy of BIMP remains stable even at very deep layers, while the energy of baseline modes does not.
In this section, we highlight the classification accuracy of BIMP on homophilic and heterophilic datasets; and the robustness of BIMP features to oversmoothing even in very deep architectures.
6.1 Classification accuracy
Homophilic datasets. We report the classification performance of our BIMP model and competitive baselines on the Cora [85], citeseer [86], Pubmed [87], CoauthorCS [88], Computers [88], and Photo [88] datasets in Table 2. We report the performance of baselines amended with various oversmoothing mitigation techniques in Appendix D.1.1. Our BIMP model outperforms all baseline and amended baseline models on all datasets.
Heterophilic datasets. We report the classification performance of our BIMP model and competitive baselines on the Texas, Wisconsin, and Cornell datasets from the CMU WebKB [89] project in Table 3. We report additional comparisons against baselines amended with various oversmoothing mitigation techniques in Appendix D.1.2. Our BIMP model outperforms all baseline and amended baseline models on all datasets.
Large graphs. We report the classification performance of BIMP and competitive baselines on the ogbn-arXiv [90] large graph dataset in Appendix D.1.3, Table 6. BIMP consistently outperforms all baseline models.
6.2 Performance at large depths
Classification accuracy. To understand the impact of depth on classification accuracy, we compare the performance of our BIMP model to baseline models over a range of network depths (hyperparameters remain fixed across depths). We report classification performance in Figure 2 (left). BIMP performs comparably to baseline models for shallow depths ($l<32$) and consistently outperforms all baseline models at greater depths. Experimental details are provided in Appendix D.2.1. We report additional comparison against discrete-depth GNNs and baselines amended with various oversmoothing mitigation techniques in Appendix D.2.1.
Dirichlet energy. To understand how oversmoothing evolves with depth, we compare the Dirichlet energy (see Equation (2)) of our BIMP model to baseline models over 1000 timesteps and plot the results in Figure 2 (right). The Dirichlet energy of our BIMP model is stable for all timesteps while the energy of baseline models diverges. Additional experimental details are provided in Appendix D.2.2. We report additional comparison against discrete-depth GNNs and baselines amended with various oversmoothing mitigation techniques in Appendix D.2.2.
7 Conclusion
In this paper, we propose an analogy between GNNs and opinion dynamics models highlighting the equivalence between oversmoothing in GNNs and opinion consensus in opinion dynamics. Through our analogy, we prove that several existing GNN algorithms are equivalent to linear opinion dynamics models which converges to consensus for all initial conditions and graph structures. Motivated by this, we introduce a novel class of continuous-depth GNNs called Behavior-inspired message passing (BIMP) GNNs which leverage the nonlinear opinion dynamics inductive bias which improves expressivity and guarantees dissensus. Experiments against competitive baselines illustrate our model’s strong performance and robustness to oversmoothing.
Limitations. The nonlinear opinion dynamics inductive bias may introduce training instabilities at larger step sizes. For example, when using the Euler method, if the step size is larger than $1/d$, where $d$ is the damping term in our dynamical update equation (see Equation (11)), each update will lead to a sign change. Since the $d$ is a hyperparameter this is not a severe limitation, and should be navigable in most settings.
Appendix A Proofs
A.1 Oversmoothing and Opinion Consensus
A.1.1 Lemma 4.1: Linear dynamics oversmooth
Any discrete-depth graph neural network with linear aggregation exhibits oversmoothing.
@@ A.1 Oversmoothing and Opinion Consensus
Without loss of generality, we consider the case where $\mathbf{A}$ is right stochastic. In this case, we can write the output of the $L$-th layer,
$\mathbf{X}^{L}=\mathbf{AX}^{L-1}\mathbf{W}^{L}=\mathbf{A}^{(L)}\mathbf{X}^{0}\mathbf{W}^{0}\cdots\mathbf{W}^{L}.$ (17)
where $\mathbf{A}^{(n)}$ denotes the $n$-th power of matrix $\mathbf{A}$.
By our GNN-OD analogy, the quantity $\mathbf{A}^{(L)}\mathbf{X}^{0}$ tends toward consensus for all initial conditions. Since this cannot be changed by a linear transformation (e.g., the product $\mathbf{W}^{0}\cdots\mathbf{W}^{L}$), $\mathbf{X}^{L}$ will also tend toward consensus. Since oversmoothing is analogous to opinion consensus, any discrete-time GNN with linear aggregation will exhibit oversmoothing for all inputs. ∎
A.1.2 Lemma 4.2: Laplacian dynamics oversmooth
Any continuous-depth graph neural network with Laplacian dynamics exhibits oversmoothing.
@@
Any GNN with Laplacian dynamics can be expressed in the form,
$\dot{\mathbf{X}}(t)=-\mathbf{DX}(t)+\mathbf{AX}(t).$ (18)
In this form, it is clear that the dynamics are equivalent to linear opinion dynamics (see Equation (6)). Since in a linear opinion dynamics model consensus is reached for all initial conditions, a GNN with Laplacian dynamics will also exhibit oversmoothing for all inputs. ∎
A.1.3 Lemma 4.3: Laplacian dynamics with an external input oversmooth
Any continuous-depth graph neural network with Laplacian dynamics and an external input will exhibit oversmoothing when the dynamics are linear.
@@
We prove Lemma 4.3 by demonstrating potential oversmoothing behavior in two popular methods: $\ell$ [48] (in Lemma A.1), GREAD-F and GREAD-FB* [49] (in Lemma A.2). The insight that the stability of linear dynamics is sensitive to the external input can be generalized to the analysis of other Laplacian dynamics with an external input. ∎
Lemma A.1 (Oversmoothing in $-\ell$).
$-\ell$ exhibits oversmoothing.